# Annotation Free Semantic Segmentation with Vision Foundation Models

### v1

@@ Abstract
Semantic Segmentation is one of the most challenging vision tasks, usually requiring large amounts of training data with expensive pixel-level annotations. With the success of foundation models and especially vision-language models, recent works attempt to achieve zero-shot semantic segmentation while requiring either large scale training or additional image/pixel-level annotations. In this work, we build a lightweight module on top of a self-supervised pretrained vision encoder to align patch features with a pre-trained text encoder. Importantly, we generate free annotations for any semantic segmentation dataset using existing foundation models and train our alignment module cost free. We use CLIP to detect objects and SAM to generate high quality object masks. Our approach can bring language-based semantics to any pre-trained vision encoder with minimal training. Our module is lightweight, uses foundation models as a sole source of supervision and shows impressive generalization capability from little training data with no annotation.
@@ 1 Introduction
In the field of computer vision, end-to-end training has long been the de facto standard, where all model parameters are optimized simultaneously for a given task or loss. Recently, large-scale and inexpensive training data has enabled the surge of foundation models [3]. These models have been conversely leveraged in order to avoid the need for both expensive annotations and extensive computational requirements, as previously required by some vision tasks [26, 19, 38, 25, 1]. While obtaining cheap annotations and leveraging pretrained models is straightforward for image classification tasks, it remains a challenge for dense prediction tasks like semantic segmentation, as existing pretrained models lack either in local feature robustness [26, 17] or in semantic awareness [38, 23, 19].
When deploying foundation models for the semantic segmentation task, there are certain properties that are desirable. First, the model has to be promptable with an open set of categories, in order to swiftly deploy it to new datasets or ontologies, without requiring extensive retraining or fine-tuning on pixel-level annotations. In this setting, we assume that there are no available pairs of image samples and pixel-wise labels. Secondly, the foundation models must generalize to different domains, and must be highly discriminative for dense recognition tasks. The open question is how to leverage existing foundation models, composing them to arrive at a unified model that covers all of the desired properties for a specific setting, and thus avoiding costly retraining.
Figure 1: Overview of our method. We propose a novel approach to annotation free semantic segmentation by composition of different foundation models. a) Label Generation: We detect possible object and background categories in an image using a pretrained image-text model (e.g. CLIP, in pink). We select image patches with high similarity to the text representation of the detected categories. Finally, we pass the location of those patches to a mask proposal network (e.g. SAM, in green). b) Visual Features Alignment: We use the generated segmentations and detected categories to align features from any arbitrary, more expressive, image encoder (e.g. DINOv2, in blue) with a pretrained text encoder (with our alignment module, in purple). c) Test-Time Inference: At test time, the newly aligned image encoder projects image features into text space. They are finally classified according to their similarity to the pre-computed vocabulary prototypes.
@@
Foundation models that align images with text descriptions present some of these properties. In particular, Contrastive language-image pretraining (CLIP [26]), trained on a large set of image-text pairs, has shown great success as a baseline for other tasks, such as zero-shot open-set image classification. It is, however, not designed for dense recognition tasks. The text-image alignment process takes the image as a whole and has no explicit notion of object localization or extent. There exists, however, some evidence that CLIP image features can be used to derive a semantic understanding of different regions within an image [43] [25], producing features that are nonetheless sub-optimal for fine dense recognition tasks such as semantic segmentation. On the other hand, recent works show that self-supervised task-agnostic pretraining methods, when applied to a sufficiently large set of unlabelled image data, can produce robust and generic visual features, suitable for any vision task. In particular, DINOv2 [25] features show generalization across different domains and are discriminative for dense recognition tasks. These models are however trained without any explicit notion of semantic meaning.
@@
In this paper, we propose a novel approach to open vocabulary semantic segmentation by composition of different foundations models as building blocks and source of supervision. The result is a model that is both generalizable, and discriminative of semantic meaning and object shape. We propose a light-weight alignment module, which projects the rich features of a self-supervised model to the embedding space of a model trained with text-image pairs. The resulting features are both highly discriminant and are grounded to semantics, while not requiring human-generated segmentation or image annotations. Our method is generic and can be applied to any pretrained image or text encoder.
Specifically, the proposed alignment module is trained in a contrastive manner to align visual features from a DINOv2 encoder with the text embedding space of a CLIP model. In order to supervise the shape and extension of objects, we generate pseudo ground truth masks from readily available images and existing dataset ontologies, by leveraging one more vision foundation model. We query the text-image model with prompts derived from a generic predefined set of labels and use the location of confident activations as prompts to Segment Anything (SAM [19]). We use the resulting masks as the supervision for the alignment module. To account for the noisy pseudo annotations, we deploy a loss function that relies on pretrained text features as anchors while optimizing the visual representations and their alignment with the text features.
@@
At test time, the alignment module projects image patches into text space, and the text-aligned patches are decoded into pixel-wise semantic segmentation features. Rather than relying on a mask proposal network, such as SAM, at test time we compute prototypes for each category of the ontology in text-feature space, and classify each output pixel according to its similarity to each prototype. This allows us to use an arbitrary category ontology to produce dataset-specific semantic segmentations on-the-fly. An overview of our method is shown in Fig. 1. Our alignment module is light and fast to train with only thousands of unannotated images and few training epochs we achieve SOTA results avoiding the need for expensive pretraining.
@@
The main contributions of this paper are: (1) We propose composing pre-trained language-image and self-supervised vision foundation models by means of a lightweight contrastive alignment module. (2) We propose a method to generate semantic segmentation masks as pseudo annotation with zero pixel or image level labels. (3) We apply the composed model to the zero-shot semantic segmentation setting, where we show generalization to never-seen-before semantic segmentation datasets. 4) We achieve state of the art results annotation-free semantic segmentation tasks.
@@ 2 Related Work
Recent advancements in large-scale pre-training have led to powerful generalist models, the so called foundation models [3], transferable to various tasks and domains. In particular, expressive vision-language models emerge when scaling contrastive pre-training on text-image pairs, as shown by CLIP [26, 17]. These models have been since applied to a range of vision problems, including open vocabulary semantic segmentation [10, 34, 34] and image generation [27]. SAM [19] is a powerful image segmentation model, trained iteratively (while providing weak annotations) on one billion images. SAM is not grounded to semantic meaning and, although there was an initial attempt by the authors to align SAM with CLIP text embeddings, no quantitative results were reported. Additionally, SAM provides segments of different granularities, making it difficult to extract object segmentations directly from a single query point.
DINO [6] and DINOv2 [25] are self-supervised image encoders that have shown to have better performance than text-aligned models in dense prediction tasks. Conversely, DINO models are not aligned with text, making their application to Open Vocabulary scenarios not straightforward. GLIP [21], and more recently GLIPv2 [39], aim to unify object detection and vision-language understanding by training on a large number of image-text pairs and bounding boxes. Grounding DINO [22] extends a closed-set detector [38] to the open-set setting by training a vision-language cross-modality decoder for cross-modality fusion. More recently, Zou et al. (2023) [44] proposed a decoder for semantic segmentation, image captioning, and visual question answering, which was pretrained on panoptic segmentation with 104K images and 4M image-text pairs. The decoder uses cross-attention between image and text queries, requiring a close-set ontology to be known in advance, thus making its application to the open-set setting limited. While other foundation models have been proposed, in this work we show an example of foundation models compositionality, where knowledge from different models (CLIP [26], SAM [19] and DINOv2 [25]) are used to provide strong open vocabulary semantic segmentation model.
@@
Open vocabulary semantic segmentation. Semantic segmentation is the task of assigning a semantic label to each pixel in a given image. In zero-shot semantic segmentation a model is expected to provide pixel-level labels for unseen categories [4]. Open vocabulary semantic segmentation can be considered as a generalized zero-shot semantic segmentation problem, one that aims at segmenting an arbitrary set of classes. Zhao et al. (2017) [40] were the first to introduce the open vocabulary setting by learning a joint embedding for pixels features and textual concepts based on WordNet.
@@
CLIP-based open vocabulary semantic segmentation. Recent works have advanced the state-of-the-art in the open vocabulary semantic segmentation problem, by leveraging self-supervised and vision-language models. Some of these approaches still require large-scale training. LSeg[20] matches the pixel embeddings to text embedding of CLIP. OpenSeg [14] learns visual-semantic alignments by aligning each word in a caption to one or few predicted masks. The general trend for open vocabulary methods is however to build on top of CLIP [8, 10, 36, 41, 28] for better generalization with less supervision. A popular line of work [35, 36, 15, 11, 37] rely on training a class-agnostic mask proposal network, leveraging CLIP embeddings for mask classification with additional techniques to strengthen CLIP image patch embeddings. These approaches still need various degrees of pixel-level supervision, which can be expensive to obtain for fine-grained categories. Training on available limited labelled datasets might thus affect their transferability and generalizability.
Open vocabulary semantic segmentation with zero pixel level supervision. One of the first approaches to zero-shot semantic segmentation with no pixel-level supervision is MaskCLIP [43] , which utilizes CLIP to generate annotations and uses them as pseudo-labels for training a segmentation network. Unlike MaskCLIP, we train a lightweight module on top of a frozen pretrained vision encoder, rather than training a complete new network. Other methods modify the CLIP encoder and update it based on a large set of image-text pairs. GroupVit [34] optimizes a hierarchical pixel grouping strategy integrated in a learned ViT model. TCL [7] trains a decoder to ground masks with language based on datasets of 12M and 3M images. Closer to our work, SAM-CLIP [32] finetunes SAM pretrained model by adding an additional CLIP head and then distilling from both CLIP and SAM using unsupervised data of 41M images. This approach tries to bring semantics to SAM but relies on large scale retraining. Our works uses SAM as a segmentation teacher and aligns a more expressive vision encoder with language, hence requiring much less training data and no model retraining. In particular, our method surpasses previous works performance using only 118 thousand unlabeled images.
@@
Training free methods. ReCo [29] builds a dataset of $K$ images for each concept from large-scale unlabelled dataset. Then a nearest neighbour approach is used to produce initial segmentation of a target concept that is then refined by DenseCLIP [28]. CLIP-DIY [33] divides an image into smaller patches that are then classified by CLIP. Patches are then aggregated and transformed to dense prediction using objectness scores from pretrained foreground-background segmentation model. In our work, we rely on SAM to provide accurate masks of detected objects. All training free methods, require a large preprocessing time and usually many steps of refinements, rendering them infeasible. We use the training-free part of our method as a source of pseudo annotations to then produce an efficient model.
Our work is the first that aligns an off the shelf pretrained and frozen vision encoder with another text encoder, at patches level with little training and light-weight alignment module; making our methods readily accessible and usable by practitioners for plug and play annotation free open vocabulary semantic segmentation.
@@ 3 Method
Semantic segmentation is the task of assigning a semantic meaning to each pixel of an image. Given the recent advancements in large pretrained models, we ask the question of whether these foundation models can be redeployed by composition to the task of semantic segmentation. We consider 3 main models. 1) CLIP [26], pretrained on a million image-text pairs, exhibiting semantic understanding of the image as a whole but not designed for object localization. 2) Purely self-supervised pretrained vision encoders, such as DINOv2 [25], producing features that transfer well to many downstream tasks, and that are consistent across similar objects and parts of objects, but without an explicit link to semantic notions. 3) SAM [19], trained to segment objects or parts of objects, but lacking proper semantic understanding. We design a method that leverages the distinctive properties of those foundation models to enable zero-shot and open vocabulary semantic segmentation.
To achieve this, we propose a two-stage approach: first, we leverage CLIP and SAM to generate pseudo semantic masks for a given vocabulary of classes. In the second stage, we use the predictions of the first stage to train a small alignment module that aligns a frozen off-the-shelf image encoder,  DINOv2 [25] with a pretrained text encoder at the patch level, resulting in a strong self-supervised semantic segmentation model.

### v2

# Fabien Despinoy

@@ Abstract
Semantic Segmentation is one of the most challenging vision tasks, usually requiring large amounts of training data with expensive pixel level annotations. With the success of foundation models and especially vision-language models, recent works attempt to achieve zeroshot semantic segmentation while requiring either large-scale training or additional image/pixel level annotations. In this work, we generate free annotations for any semantic segmentation dataset using existing foundation models. We use CLIP to detect objects and SAM to generate high quality object masks. Next, we build a lightweight module on top of a self-supervised vision encoder, DinoV2, to align the patch features with a pretrained text encoder for zeroshot semantic segmentation. Our approach can bring language-based semantics to any pretrained vision encoder with minimal training. Our module is lightweight, uses foundation models as the sole source of supervision and shows impressive generalization capability from little training data with no annotation.
@@ 1 Introduction
Large-scale and inexpensive training data has recently enabled the surge of foundation models in computer vision [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al.]. These models have been employed to avoid expensive annotations and computational requirements of many vision tasks [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al., Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al., Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al.]. Leveraging foundation models for semantic segmentation requires the model to produce pixel-wise predictions generalizable to new datasets, domains and ontologies. Therefore, the model must be 1) promptable with an open set of categories and 2) highly discriminative for dense recognition tasks. Consequently, deploying foundation models to obtain cheap annotation for semantic segmentation is challenging as existing models lack either semantic awareness [Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie, Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] or local feature robustness [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Jia et al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and Duerig].
[…]
@@
In particular, there exists some evidence that Contrastive Language-Image Pretraining (CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.]), trained on a large set of image-text pairs, can be used to derive a semantic understanding of different regions within an image [Zhou et al.(2022)Zhou, Loy, and Dai, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.]. However, CLIP features are sub-optimal for dense recognition tasks such as semantic segmentation. On the other hand, self-supervised task-agnostic pretraining methods applied to a sufficiently large set of unlabelled image data have shown great success in producing robust and generic visual features, suitable for any vision task. Specifically, DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] features show generalization across different domains and are discriminative for dense recognition tasks. DINO models are however trained without any explicit notion of semantic meaning  [Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.].
@@
In this paper, we propose a novel approach to open vocabulary semantic segmentation by composition of different foundations models as building blocks and source of supervision. The result is a model that is highly discriminant, generalizable and grounded to semantic meaning and object shape without requiring human-generated segmentation or image annotations. We propose a lightweight alignment module which is trained in a contrastive manner to align visual features from a DINOv2 encoder with the text embedding space of a CLIP model. In order to supervise the shape and extension of objects, we generate pseudo ground truth masks from readily available images and existing dataset ontologies, by leveraging one more vision foundation model, Segment Anything (SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]).
@@
At test time, the alignment module projects image patches into text space, and the text-aligned patches are decoded into pixel-wise semantic segmentation features. Each pixel is classified according to its similarity to the text embeddings of a target ontology. This enables the model to use any ontology to produce dataset-specific semantic segmentations on-the-fly. An overview of our method coined as FMbSeg is shown in Fig. 1. Our alignment module is light and fast and with only thousands of unannotated images and few training epochs we achieve SOTA results avoiding the need for expensive pretraining.
@@
The main contributions of this paper are: (1) We propose a method to generate semantic segmentation pseudo annotations with zero pixel or image level labels using CLIP and SAM. (2) We propose composing pretrained language-image and self-supervised vision foundation models by means of a lightweight contrastive alignment module. (3) We apply the composed model to the zeroshot semantic segmentation setting, where we show generalization to never-seen-before semantic segmentation datasets. 4) We achieve state of the art results on annotation-free semantic segmentation task.
@@ 2 Related Work
Recent advancements in large-scale pretraining have led to powerful generalist models, transferable to various tasks and domains  [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al.]. In particular, expressive vision-language models emerge when scaling contrastive pretraining on text-image pairs, as shown by ALIGN and CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Jia et al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and Duerig]. These models have been since applied to a range of vision problems, including open vocabulary semantic segmentation [Ding et al.(2022)Ding, Xue, Xia, and Dai, Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] and image generation [Ramesh et al.(2021)Ramesh, Pavlov, Goh, Gray, Voss, Radford, Chen, and Sutskever]. SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] is a powerful image segmentation model, trained iteratively with weak annotations on one billion images. SAM is not grounded to semantic meaning and although there was an initial attempt by the authors to align SAM with CLIP text embeddings, no quantitative results were reported. DINO [Caron et al.(2021)Caron, Touvron, Misra, Jégou, Mairal, Bojanowski, and Joulin] and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] are self-supervised image encoders demonstrating high performance in dense prediction tasks. However, DINO models are not aligned with text, making their application to Open Vocabulary scenarios not straightforward. In this work we show an example of foundation models compositionality, where knowledge from different models (CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.]) are used to provide strong open vocabulary semantic segmentation model.
@@
Open vocabulary semantic segmentation. Zeroshot semantic segmentation requires a model to provide pixel level labels for unseen categories [Bucher et al.(2019)Bucher, Vu, Cord, and Pérez]. Open vocabulary semantic segmentation generalizes this problem aiming at segmenting an arbitrary set of classes. Zhao et al. (2017) [Zhao et al.(2017)Zhao, Puig, Zhou, Fidler, and Torralba] were the first to introduce the open vocabulary setting by learning a joint embedding for pixels and textual concepts based on WordNet.
@@
CLIP-based methods. The general trend for open vocabulary methods is to build on top of CLIP for better generalization with less supervision [Li et al.(2022b)Li, Zhang, Zhang, Yang, Li, Zhong, Wang, Yuan, Zhang, Hwang, et al., Zhang et al.(2022b)Zhang, Zhang, Hu, Chen, Li, Dai, Wang, Yuan, Hwang, and Gao, Chen et al.(2023)Chen, Li, Lim, Torralba, and Zhao, Ding et al.(2022)Ding, Xue, Xia, and Dai, Xu et al.(2023)Xu, Xiong, Ding, and Tu, Zhong et al.(2022)Zhong, Yang, Zhang, Li, Codella, Li, Zhou, Dai, Yuan, Li, et al., Rao et al.(2022)Rao, Zhao, Chen, Tang, Zhu, Huang, Zhou, and Lu]. LSeg[Li et al.(2022a)Li, Weinberger, Belongie, Koltun, and Ranftl] matches the pixel embeddings to text embedding of CLIP. OpenSeg [Ghiasi et al.(2022)Ghiasi, Gu, Cui, and Lin] learns visual-semantic alignments by aligning each word in a caption to one or few predicted masks. A popular line of work rely on training a class-agnostic mask proposal network, leveraging CLIP embeddings for mask classification with additional techniques to strengthen CLIP image patch embeddings [Xu et al.(2022b)Xu, Zhang, Wei, Lin, Cao, Hu, and Bai, Xu et al.(2023)Xu, Xiong, Ding, and Tu, Han et al.(2023)Han, Zhong, Li, Han, and Ma, Ding et al.(2023)Ding, Wang, and Tu, Yu et al.(2023)Yu, He, Deng, Shen, and Chen]. These works still require a various degrees of pixel-level supervision, which can be expensive to obtain for fine-grained categories. Training on limited available labelled datasets might thus affect their generalization.
@@
Training-free methods. ReCo [Shin et al.(2022)Shin, Xie, and Albanie] builds a dataset of $K$ images for each concept from large-scale unlabelled dataset. Then a nearest neighbour approach is used to produce initial segmentation of a target concept that is then refined by DenseCLIP [Rao et al.(2022)Rao, Zhao, Chen, Tang, Zhu, Huang, Zhou, and Lu]. CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni] divides an image into smaller patches that are then classified by CLIP. Patches are then aggregated and transformed to dense prediction using objectness scores from pretrained foreground-background segmentation model. In our work, we rely on SAM to provide accurate masks of detected objects. All training free methods, require a large preprocessing time and usually many steps of refinement, rendering them infeasible. We use the training-free part of our method (section 3.2) as a source of pseudo annotations to produce an efficient model.
Training-based methods without pixel level supervision. Many methods modify the CLIP encoder and update it based on a large set of image-text annotated pairs. GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] optimizes a hierarchical pixel grouping strategy integrated in a learned ViT model. TCL [Cha et al.(2023)Cha, Mun, and Roh] trains a decoder to ground masks with language based on datasets of 12M and 3M images. Closer to our work, MaskCLIP [Zhou et al.(2022)Zhou, Loy, and Dai] utilizes CLIP to generate annotations for training a complete segmentation network from scratch. Unlike this work, we train a lightweight module on top of a frozen pretrained vision encoder. SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] brings semantics to SAM by large-scale fine-tuning and distilation from both SAM and CLIP models with 41M unannotated images. Our work employs SAM as a segmentation teacher and aligns a more expressive vision encoder with CLIP’s pretrained language encoder, hence requiring much less training data and no model retraining. Particularly, our method surpasses previous works performance using only 118 thousand unlabeled images. Our work is the first that aligns off the shelf vision and text encoders, at patches level with little training and a lightweight alignment module; making our methods readily accessible and usable by practitioners for plug and play annotation-free open vocabulary semantic segmentation.
@@ 3 Method
In this work, we deploy 3 foundation models by composition for the task of semantic segmentation: 1) CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], pretrained on a million image-text pairs, exhibiting semantic understanding of the image as a whole but not designed for object localization. 2) SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.], trained to segment objects or parts of objects, but lacking proper semantic understanding. 3) DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], producing features that transfer well to many downstream tasks, and that are consistent across similar objects and parts of objects, but without an explicit link to semantic notions. We design a method that leverages the distinctive properties of those foundation models to enable zeroshot and open vocabulary semantic segmentation.
To achieve this, we propose a two-stage approach: first, we leverage CLIP and SAM to generate pseudo semantic masks for a given vocabulary of classes. In the second stage, we use the predictions of the first stage to train a small alignment module that aligns a frozen off-the-shelf image encoder,  DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], with a pretrained text encoder at the patch level, resulting in a strong self-supervised semantic segmentation model.